OpenStudy (anonymous):

is the degree of a4b-7a2=4 ?

OpenStudy (aroub):

no..

OpenStudy (anonymous):

ok will you please teach me hoe to get it right:)

OpenStudy (aroub):

sure..

OpenStudy (anonymous):

thanks i meant how

OpenStudy (anonymous):

sorry i put the wrong problem the question is , is the degree a4b-7a2b2=4

OpenStudy (aroub):

in this specific question u have 2 terms in every term u have to add the exponents like for example the first one a^4b u have to add 4+1 =5 ( i guess u forgot to add the 1 right?) and for the second term 7a2 its only 2 so which oe is larger that's the degree.. lol sorry i know im not that good at explaining... oh okay well do the same thing why don't u try this time??

OpenStudy (anonymous):

 so degree=5